OpenStudy (anonymous):

Find the surface area of a sphere with a circumference of 40 ft. Round to the nearest tenth. (1 point)

OpenStudy (anonymous):

Surface are of a sphere is given by : 4 times pi times radius squared = 4xpix(40ft) squared

OpenStudy (anonymous):

This will equal around 6400pi ft squated

OpenStudy (anonymous):

=20106.2 ft2

OpenStudy (anonymous):

not a option i have

OpenStudy (anonymous):

6.4 509.3 245.6 40.5 these are my options i have

OpenStudy (anonymous):

@shubhamk u there

OpenStudy (anonymous):

What are the units mentioned in the answer

OpenStudy (anonymous):

I was a lot wrong circumference is given by 2times pi times radius=40 So. radius =20/pi Place this in formula for surface area given. by :4 x pi x radius squared = 1600 / pi = 509.3 T